6. a) You want to amplify an input signal with 10 mV amplitude and 50 MHz frequency. 

Which characteristics of op-amp will be significant in this case? What will be the 

value of these parameters? Why?

b) Design and draw a circuit using an op-amp to get + 8 V at output for input less than 

+4 V and -8 V at output for input greater than +4 V. 


1
Expert's answer
2021-04-28T07:43:57-0400

To be given in question

Voltage (v)=10mv

Frequency (f)=50Hz

To be asked in question

(a) slew rate

(b) circuit design


We know that

(a) Slew rate (SR)= "\frac{dv}{dt}"

Slew rate :- output voltage change with respect to time this is called slew rate

Slew rate ="10\times10^{-3}\times50\times10^6 volt/sec"

SR ="5\times10^4meter/sec"

(b) when Vo=8volt



V"_{i}=<4volt"

Gain is "\beta=\frac{V_{o}}{V_{i}}"

"\beta>2"

"V_{0}=-8volt"

"V_{i}=4volt"

"\beta=\frac{-8}{4}=-2"

"\beta<2"

Condition follow